# Institutionalizing Folk Theories of Algorithms: How Multi-Channel Networks (MCNs) Govern Algorithmic Labor in Chinese Live-Streaming Industry

### v1

@@ 4.2. External Folk Theories: Simplified Narrative to Manage and Govern Streamers
In this section, we analyze how MCNs construct and disseminate folk theories of algorithms to live-streamers. Unlike the probabilistic and speculative nature of internal folk theories held by MCN staff, these outward-facing theories of algorithms targeting streamers are packaged as confident, simplified rules that make the algorithm seem legible and controllable. We identify two primary ways in which these theories are strategically deployed: (1) by translating algorithmic logic into concrete, observable metrics such as appearance and equipment quality (subsubsection 4.2.1); (2) by framing algorithmic success as attainable through discipline, investment, and compliance (subsubsection 4.2.2). These theories do not merely inform behavior that shape streamers’ beliefs, extract their investment, and sustain a system in which hope becomes both affective infrastructure and a mechanism of labor control.
@@ 4.2.1. Simplifying Algorithmic Logic Through Observable Metrics
In the face of algorithmic opacity and instability, MCNs firstly offer simplified but confident explanations to live-streamers that translate the platform’s black-box logic into concrete, observable metrics. Without access to the actual workings of recommendation systems, MCNs construct a seemingly actionable framework for live-stramers grounded in visible variables, such as physical appearance, equipment quality, and early user engagement, to make the algorithm appear understandable and responsive to effort.
For example, M2 usually explains to new streamers that physical appearance plays a decisive role in gaining algorithmic exposure by increasing early viewer retention: “The audience will only give a live-streamer three seconds to show themselves and keep them. When a live-streamer is particularly attractive or has a humorous appeal, they are more likely to gain audience favor. This results in higher viewer retention and influx rates, leading the recommendation algorithm to classify them as popular live-streamers and, consequently, grant them more exposure.” This logic, while unverified, is widely internalized by streamers.
Another common metric emphasized by MCNs is the quality of equipment. M4 explained that many platforms set technical thresholds for stream quality, and that insufficient equipment could lead to streams being algorithmically filtered out: “If your sound and visuals are poor, people drop out quickly, and that hurts your retention rate. Algorithms pick up on that. So we always recommend at least the basics, e.g., lights, microphones, stable connection.” Streamers echoed this perceived necessity. L5 noted that even renting equipment was necessary for basic visibility:“Whether renting or buying equipment, it at least gives you a chance to gain some recommended followers and earn a bit of money. Without investing in equipment, it’s nearly impossible to be visible to viewers.” As shown in Figure 1, MCNs distribute tiered technical audio equipment in ways that symbolically link production quality with algorithmic visibility, reinforcing an externally imposed folk theory that equates financial investment in equipment with success.
@@
What often begins as a recommendation to improve a single technical element, such as audio clarity or lighting, gradually expands into a broader vision of the “ideal” streaming setup. MCNs do not simply promote individual tools; they encourage streamers to view the entire streaming environment as a form of algorithmic currency. Ring lights, DSLR cameras, soundproofed backdrops, multi-angle mounts, branded filters, and LED mood lighting are framed as signs of professionalism and algorithm-readiness. As streamers internalize this logic, the line between technical sufficiency and aesthetic perfection becomes blurred. A basic setup is no longer seen as enough to ”meet the algorithm’s baseline”; instead, streamers feel pressured to constantly upgrade in order to stay competitive within an ever-rising standard.
However, most MCN staff themselves acknowledged the limitations of this external theory for streamers (which they don’t believe internally). M9 pointed out the asymmetry between perceived causes of success and actual algorithmic outcomes: “We can assume that average looks or limited equipment make it difficult to receive algorithmic recommendations. However, this assumption cannot be reversed. We can’t conclude that being attractive or having better equipment will guarantee algorithmic recommendations. Our company’s data doesn’t support this.”
Despite this internal skepticism, MCNs continue to promote the idea to live-streamers that algorithmic success is predictable and dependent on specific surface-level traits. They construct the algorithm as a controllable input-output system: if streamers follow the right steps, if maintain a certain visual style, upgrade equipment, stream for long enough, and use the right tags, they are told they will be rewarded with visibility and growth. This framing simplifies algorithmic complexity into a checklist of behaviors and attributes, turning uncertainty into a set of seemingly manageable variables.
Because this theory of algorithm constructed by MCNs feels actionable, most streamers often accept it without question. They internalize the idea that if they fail to gain traction, it is because one of these variables was insufficient. Perhaps their lighting was not professional enough, their background not aesthetic enough, their streaming duration too short, or their performance too flat. This belief system allows MCNs to continuously raise performance benchmarks, always pointing to another metric that could be optimized. For streamers, each failed attempt becomes a reason to try harder, upgrade further, or self-correct, rather than question the underlying opacity or instability of the algorithm itself. In this way, MCNs maintain authority while pushing laborers into a cycle of self-blame and incremental reinvestment. Crucially, this theory also creates a steady revenue stream for MCNs, as streamers are encouraged to repeatedly invest—purchasing equipment, renting better spaces, or enrolling in paid training sessions—all under the belief that these expenditures will enhance their algorithmic visibility and lead to success.
@@ 4.2.2. Framing Algorithmic Success as Attainable and Controllable
A second common way MCNs explain the algorithm is by framing it as attainable and controllable, a fair system that rewards effort, discipline, and the right technical setup. Despite the algorithm’s underlying opacity and volatility, MCNs portray it as a rational mechanism that can be “understood” and “tamed” through compliance with certain behavioral and technical standards. This framing offers streamers a sense of structure and hope, suggesting that visibility is within reach if they follow instructions and persist.
L3 described how this theory of algorithms shaped her initial investment in equipment: “Those who can’t afford live-streaming equipment have to rely on income from other sources to make an initial investment for the equipment [if they believe live-streaming could earn more money].”
@@
This theory is not just implicit. It is actively constructed by MCNs during recruitment and onboarding. New streamers are told that the algorithm can be “understood” and “tamed,” and that their success depends on how closely they adhere to MCN advice and technical standards. This framing channels attention away from platform bias or structural inequality, and toward individual action and responsibility.
@@
Crucially, MCNs often present live-streaming platforms and their algorithmic infrastructures as less discriminatory and more open than other forms of labor or public-facing institutions. In particular, our sampled MCNs in China emphasize that streamers from rural, migrant, or female backgrounds may find more opportunities online than in traditional workplaces. Unlike offline industries where class, gender, accent, or education may limit mobility, MCNs describe algorithmic systems as blind to such characteristics, privileging performance, persistence, and technique instead. Several streamer interviewees (n=7) recalled being told during onboarding sessions that “the algorithm doesn’t care who you are, only what you do”, a message that resonated deeply with streamers from marginalized backgrounds.
@@
This perception helps explain why live-streaming, despite its precarity, continues to attract disproportionately high numbers of rural migrants and young women. The promise of algorithmic neutrality, reinforced by MCNs’ messaging, offers both a psychological appeal and a perceived escape from entrenched discrimination elsewhere. By portraying the algorithm as neutral and open to anyone, MCNs produce not just training routines for live-streamers, but an aspirational and moral framework in which effort is rewarded, and structural inequalities appear bypassed.
However, as most of MCNs interviewees (N=8) agreed, the actual support that MCNs provide is often minimal. They may offer basic resources, such as startup equipment, help with platform registration, and initial streaming advice, but rarely offer deeper guidance about how to gain algorithmic visibility. What matters more is the theories of algorithms they construct: that the algorithm is fair, that it can be tamed, and that success is attainable through individual effort. This story is powerful not because it is empirically accurate, but because it motivates.
For MCNs, this external version about theory of algorithms serves a clear purpose: it attracts a continuous influx of new streamers. Even if only a small percentage achieve visibility, the system is sustained by the hope of success and the belief that the algorithm rewards discipline. For streamers, especially those from marginalized backgrounds with few algorithm literacy, this external folk theory provides a rare sense of possibility, even if the odds remain opaque and the outcomes unpredictable.
@@
Crucially, it is precisely this belief that the algorithm can be learned, tamed, and ultimately trusted sustains repeated investment from streamers. Streamers are told that success is not arbitrary, but a matter of strategy, effort, and consistency. The algorithm, unlike human employers or traditional gatekeepers, is portrayed as objective and responsive by MCNs. For many streamers, especially those excluded from stable or upwardly mobile labor markets, the promise of algorithmic fairness feels both empowering and worth pursuing. This perception justifies the continued purchase of equipment, the adoption of performance routines, and the sacrifice of time and emotional energy. Even when results fall short, the theories remains intact: if one stream fails, the next might succeed. In this way, MCNs not only mobilize belief—they build a system in which belief becomes labor, and hope becomes infrastructure.
5. Findings: Institutionalizing Folk Theories: How MCNs Operationalize Algorithmic Beliefs for Labor Management (RQ2)
@@
Building on section 4, which examined how MCNs construct internal and external folk theories of algorithms, we now turn to how these interpretations are institutionalized into labor control. First, MCNs embed algorithmic expectations into everyday routines, through onboarding, coaching, and evaluations, repeating the message that effort, emotion, and consistency will be rewarded by the algorithm. This moralized framing aligns with the external folk theory that success is earned, motivating streamers to self-discipline and persist. At the same time, MCNs internally acknowledge the algorithm’s volatility and shift the burden of failure onto streamers through this moral frame, who blame themselves rather than systemic opacity (subsection 5.1). Second, MCNs formalize these beliefs through structured schedules and aesthetic standards, reinforcing the idea that algorithmic success can be optimized. Streamers are encouraged to upgrade equipment, refine routines, and purchase services—all justified by the external folk theory. Yet MCNs, guided by internal theories, understand the algorithm offers no guarantees. Their profits rely not on streamer success, but on sustaining belief and dependence on paid services (subsection 5.2).
@@ 5.1. Transforming Platform Speculation into Moral Labor Expectations
As discussed in the previous subsection 4.2, MCNs interpret opaque and unstable platform algorithms by translating them into simplified rules and actionable advice. These interpretations are not neutral; they serve specific organizational interests by creating the illusion that algorithmic success can be achieved through discipline, effort, and optimization. In this section, we show how such interpretations do not simply circulate as loose theories. They are systematically leveraged to shape how streamers work, behave, and evaluate themselves.
@@
One major way this happens is through the transformation of speculative ideas about the algorithm into moral labor expectations. Beliefs such as “the algorithm rewards long hours” or “visual quality guarantees recommendations” are framed not as possibilities, but as certainties. These claims are repeatedly emphasized in onboarding materials, group chat guidance, and informal coaching, where they gain authority as commonsense truths. Over time, they become internalized by streamers as standards of professional conduct.
@@
This process marks the point where algorithmic interpretation becomes algorithmic governance. Instead of merely helping streamers understand how the platform might work, MCNs use these beliefs to define how streamers should behave. The burden of visibility is shifted entirely onto the individual: if streamers are not being recommended, it is because they are not working hard enough, not investing enough, or not attractive or expressive enough. L6 recalled: “The MCNs I contacted told me I only needed to stream every day. Just show up, and the algorithm will recognize you. I thought, okay, then it’s just about effort.”
@@
This framing naturalizes structural opacity as personal failure. M2 explained: “Most streamers think it’s their fault. That’s what keeps them going. They keep chasing that moment when the algorithm ‘notices’ them.” Even when outcomes do not improve, many streamers continue to blame themselves. L7 described her own response: “They said if I streamed every day and did more facial expressions, I’d get algorithmic recommendations. So I smile even when I’m exhausted, because maybe today is the day the algorithm sees me.”
@@
These internalized expectations stem directly from the way MCNs frame the algorithm in the first place for streamers, as something knowable, responsive, and fair. Because streamers believe the system rewards the right behavior, they are willing to endure exhaustion, self-doubt, and financial risk. This makes the MCN’s job easier. As M11 put it: “We MCNs don’t need to spend much. If they streamers believe in the algorithm, they’ll work hard for free. That’s the best-case scenario for any company.”
Streamers are not being forced to work harder; they are being persuaded that working harder is their own path to success. MCNs reinforce this belief system by presenting guidance not as command, but as mentorship. L18 explained: “My manager never says I must do something. She just says, ‘Other streamers who do X or Y tend to get recommended more.’ Then she might show me screenshots. It’s never a command, but I always feel like if I don’t follow, I’ll miss out.”
@@
What appears as care or mentorship is, in fact, a form of indirect control. Streamers act based on what they believe the algorithm expects, but those beliefs are shaped by the MCN’s theories. In L22’s case, her makeup choices were labeled as insufficient by MCNs not because of any clear platform rule, but because the MCN had already instilled the idea that visual performance is central to algorithmic success. This claim itself emerges from the broader uncertainty of algorithmic systems: in the absence of transparent criteria, MCNs gain interpretive authority, turning speculation into behavioral norms. As L22 put it, “They didn’t force me, but I felt like if I didn’t improve my makeup, I was being lazy or unprofessional. It felt like a personal failing.”
@@ 5.2. Embedding Algorithmic Beliefs into Material and Contractual Structures
The interpretive work described in the previous section does not stop at narrative or motivation. MCNs extend their algorithmic explanations into the material, aesthetic, and contractual arrangements that define how streamers work. What begins as speculative beliefs, such as the idea that visual clarity or consistent appearance leads to higher algorithmic visibility, is gradually institutionalized into binding routines and financial commitments.
@@
A central mechanism is the pressured acquisition of high-end equipment. MCNs frequently tell streamers that clearer visuals and smoother audio will improve their ranking, reinforcing this belief through formal or informal expectations. L19 described her experience: “I spent 10,000 RMB on a Sony camera because the MCN said it would help me algorithmic visibility rank better. But nothing changed. They kept saying, maybe I need to improve my makeup or get a new background.”
@@
This expectation extends far beyond cameras. Ring lights, microphones, mixers, reflective backdrops, and LED panels are all framed as essential to “meeting the algorithm’s baseline.” In some cases, MCNs distribute internal “technical evaluation reports” that nudge streamers toward upgrades. M12 explained how this theory benefits the company: “Most streamers don’t buy equipment all at once. They rent from us first. If they start earning, they’ll invest more. If they don’t, at least we make money from the rentals.”
@@
Alongside hardware, MCNs standardize streamers’ aesthetic presentation. Guidelines often dictate not only tone and posture, but makeup, camera angles, and even clothing schedules. M9 described a typical protocol: “They’re not allowed to wear the same outfit in a two-week cycle. We even advise which colors look better on camera. Some of them rent clothes from us. We keep it on their account.”
@@
These requirements are enforced not just rhetorically, but contractually through studio rules, performance evaluations, and surveillance systems. Even streamers who work from home are monitored. M4 explained: “We ask them to share their live link with a manager who checks if the lighting, camera, background are aligned with expectations. If not, they get a warning. If the setup is too basic, we suggest they pause streaming until they upgrade.”
@@
In some cases, MCNs also manage the physical layout of labor. Many operate shared studio spaces where streaming rooms are tiered by performance level. Higher-ranked streamers receive better infrastructure and services. L16 described how space and performance were linked: “At first, I shared a basement room with three other girls. It was tiny, and the internet cut out often. When I ranked higher one month, I moved to a private room upstairs. But I had to pay more for it, nearly 3,000 RMB a month.” These layered expectations create cycles of escalating cost and performance pressure. L20 recalled: “I spent most of my early earnings on room rental, clothes, gear. I didn’t even realize I was in debt until month three. But by then, I didn’t dare quit; I was too invested.” Streamers are often told that these are necessary “investments” in their future algorithmic visibility. But the platform algorithm offers no guarantees. L22 shared her experience: “I borrowed money from my cousin to redecorate my stream room. I told myself, if I looked more professional, I’d get better exposure. But nothing happened. Now I just feel stuck.” As shown in  Figure 2, the material differentiation of studio spaces reflects a managed hierarchy of visibility and professionalism, reinforcing MCNs’ externally imposed belief that algorithmic success is contingent on aesthetic and technical upgrades.
@@
As illustrated in subsubsection 4.2.1, MCNs build streamers’ belief that algorithmic success depends on visible performance cues and technical polish. Once internalized, these beliefs become a foundation for control: they justify expenditures, reinforce discipline, and make streamers responsible for conditions over which they have limited influence. M12 captured the disciplining effect of this folk algorithmic theory: “We don’t force anyone to stay. But when you’ve rented equipment, decorated your room, signed the platform contract; it’s hard to walk away. You want to try a bit longer, just in case.”
For MCNs, this internalization is not just a behavioral outcome—it is a business model. Once streamers adopt the belief that algorithmic success can be engineered through effort and appearance, they become willing customers for MCN-provided goods and services: studio rentals, training sessions, lighting kits, camera upgrades, and branding packages. These purchases are not simply one-time decisions, they form an ongoing cycle of reinvestment, driven by the hope that one more improvement will finally unlock algorithmic visibility.

### v2

@@ 4.2. External Interpretation: Simplified Narrative to Manage and Govern Streamers
Unlike the probabilistic and speculative nature of internal algorithmic interpretations held by MCN staff, these outward-facing interpretations of algorithms targeting streamers are packaged as confident, simplified rules that make the algorithm seem legible and controllable. We identify two primary ways in which these algorithmic interpretations are strategically deployed: (1) by translating algorithmic logic into concrete, observable metrics such as appearance and equipment quality (subsubsection 4.2.1); (2) by framing algorithmic success as attainable through discipline, investment, and compliance (subsubsection 4.2.2). Here we focus more on the content of these beliefs and we will further analyze why MCNs constructed these interpretations in such ways for labor management in section 5.
@@ 4.2.1. Simplifying Algorithmic Logic Through Observable Metrics
In the face of algorithmic opacity and instability, MCNs firstly offer simplified but confident explanations to live-streamers that translate the platform’s black-box logic into concrete, observable metrics. For example, M2 usually explains to new streamers that physical appearance plays a decisive role in gaining algorithmic exposure by increasing early viewer retention: “The audience will only give a live-streamer three seconds to show themselves and keep them. When a live-streamer is particularly attractive, they are more likely to gain audience favor. This results in higher viewer retention and influx rates, leading the recommendation algorithm to classify them as popular live-streamers and, consequently, grant them more exposure.” This logic, while unverified, is widely internalized by streamers.
Another common metric emphasized by MCNs is the quality of equipment. M4 explained that many platforms set technical thresholds for stream quality, and that insufficient equipment could lead to streams being algorithmically filtered out: “If your sound and visuals are poor, people drop out quickly, and that hurts your retention rate. Algorithms pick up on that. So we always recommend at least the basics, e.g., lights, microphones, stable connection.” Streamers echoed this perceived necessity. L5 noted that even renting equipment was necessary for basic visibility:“Whether renting or buying equipment, it at least gives you a chance to gain some recommended followers and earn a bit of money. Without investing in equipment, it’s nearly impossible to be visible to viewers.” As shown in Figure 1, MCNs distribute tiered technical audio equipment in ways that symbolically link production quality with algorithmic visibility, reinforcing an externally imposed algorithmic interpretations that equates financial investment in equipment with success.
@@
What often begins as a recommendation to improve a single technical element, such as audio clarity or lighting, gradually expands into a broader vision of the “ideal” streaming setup. MCNs do not simply promote individual tools; they encourage streamers to view the entire streaming environment as a form of algorithmic currency. Ring lights, DSLR cameras, soundproofed backdrops, multi-angle mounts, branded filters, and LED mood lighting are framed as signs of professionalism and algorithm-readiness. As streamers internalize this logic, the line between technical sufficiency and aesthetic perfection becomes blurred. A basic setup is no longer seen as enough to “meet the algorithm’s baseline” (M9); instead, streamers feel pressured to constantly upgrade in order to stay competitive within an ever-rising standard.
MCNs continue to construct the algorithm as a controllable input-output system: if streamers follow the right steps, if maintain a certain visual style, upgrade equipment, stream for long enough, and use the right tags, they are told they will be rewarded with visibility and growth. This framing simplifies algorithmic complexity into a checklist of behaviors and attributes.
Because this interpretation of algorithm constructed by MCNs feels actionable, most streamers often accept it without question. They internalize the idea that if they fail to gain traction, it is because one of these variables was insufficient. Perhaps their lighting was not professional enough, their background not aesthetic enough, their streaming duration too short, or their performance too flat. This belief system allows MCNs to continuously raise performance benchmarks, always pointing to another metric that could be optimized.
@@ 4.2.2. Framing Algorithmic Success as Attainable and Controllable
A second common way MCNs explain the algorithm is by framing it as attainable and controllable, a fair system that rewards effort, discipline, and the right technical setup. This interpretation is actively constructed by MCNs during recruitment and onboarding. New streamers are told that the algorithm can be tamed, and that their success depends on how closely they adhere to MCN advice and technical standards. This framing channels attention away from platform bias or structural inequality, and toward individual action and responsibility.
@@
MCNs often present live-streaming platforms and their algorithmic infrastructures as less discriminatory and more open than other forms of labor or public-facing institutions. In particular, our sampled MCNs in China emphasize that streamers from rural, migrant, or female backgrounds may find more opportunities online than in traditional workplaces. Unlike offline industries where class, gender, accent, or education may limit mobility, MCNs describe algorithmic systems as blind to such characteristics, privileging performance, persistence, and technique instead. Several streamer interviewees (n=7) recalled being told during on-boarding sessions that “the algorithm doesn’t care who you are; everyone’s talents deserve to be seen”, a message that resonated deeply with streamers from marginalized backgrounds.
@@
This perception helps explain why live-streaming, despite its precarity, continues to attract disproportionately high numbers of rural migrants and young adults, especially women. The promise of algorithmic neutrality, reinforced by MCNs’ messaging, offers both a psychological appeal and a perceived escape from entrenched discrimination elsewhere.
@@
It is precisely this belief that the algorithm can be tamed sustains repeated investment from streamers. Streamers are told that success is not arbitrary, but a matter of effort. The algorithm, unlike human employers or traditional gatekeepers, is portrayed as objective and responsive by MCNs. For many streamers, especially those excluded from stable or upwardly mobile labor markets, the promise of algorithmic fairness feels both empowering and worth pursuing. This perception justifies the continued purchase of equipment, the adoption of performance routines, and the sacrifice of time and emotional energy. Even when results fall short, the interpretations remains intact: if one stream fails, the next might succeed.
5. Findings: Labor Management as a Result of MCNs’ Discursive Construction of Algorithms (RQ2)
@@
Building on section 4, which examined how MCNs construct internal and external interpretation of algorithms, we now turn to how these interpretations are institutionalized into labor control. First, MCNs embed algorithmic expectations into everyday routines, through on-boarding, coaching, and evaluations, repeating the message that effort, affect, and consistency will be rewarded by the algorithm. This moralized framing aligns with the external interpretation that success is earned, motivating streamers to self-discipline and persist. At the same time, MCNs internally acknowledge the algorithm’s volatility and shift the burden of failure onto streamers through this moral frame, who blame themselves rather than systemic opacity (subsection 5.1). Second, MCNs formalize these beliefs through structured schedules and aesthetic standards, reinforcing the idea that algorithmic success can be optimized. Streamers are encouraged to upgrade equipment, refine routines, and purchase services—all justified by the external interpretation. Yet MCNs, guided by internal interpretation, understand the algorithm offers no guarantees. Their profits rely not on streamer success, but on sustaining belief and dependence on paid services (subsection 5.2).
@@ 5.1. Transforming Platform Speculation into Moral Labor Expectations
As discussed in the previous subsection 4.2, MCNs interpret opaque and unstable platform algorithms by translating them into simplified rules and actionable advice. These interpretations serve specific organizational interests by creating the illusion that algorithmic success can be achieved through discipline, effort, and optimization. In this section, we show how such interpretations are systematically leveraged into labor management to shape how streamers work and evaluate themselves.
@@
One major way this happens is through the transformation of speculative ideas about the algorithm into moral labor expectations. Beliefs such as “the algorithm rewards long hours” or “visual quality guarantees recommendations” are framed as certainties. These claims are repeatedly emphasized in on-boarding materials, group chat guidance, and informal coaching, where they gain authority as commonsense truths. Over time, they become internalized by streamers as standards of professional conduct.
@@
Instead of merely helping streamers understand how the platform might work, MCNs use these beliefs to define how streamers should behave. The burden of visibility is shifted entirely onto the individual: if streamers are not being recommended, it is because they are not working hard enough, not investing enough on equipments, or not physically attractive or expressive enough, i.e., transforming the responsibility onto individual efforts. L6 recalled: “The MCNs I contacted told me I only needed to stream every day. Just show up, and the algorithm will recognize you. I thought, okay, then it’s just about effort.”
@@
This framing naturalizes structural opacity as personal failure. M2 explained: “Most streamers think it’s their fault. That’s what keeps them going. They keep chasing that moment when the algorithm ‘notices’ them.” Even when outcomes do not improve, many streamers continue to blame themselves. L7 described his own response: “They said if I streamed every day with more positive emotions expressions, I’d get algorithmic recommendations. So I smile even when I’m exhausted, because maybe today is the day the algorithm sees me.”
@@
Because streamers believe the system rewards the right behavior, they are willing to endure exhaustion, self-doubt, and financial risk. This makes the MCN’s job easier. As M11 put it: “We MCNs don’t need to spend much. If they streamers believe in the algorithm, they’ll work hard for free. No other incentives or strict management are needed. That’s the best-case scenario for any company.”
@@
Streamers act based on what they believe the algorithm expects, but those beliefs are largely shaped by the MCN’s interpretation. In L22’s case, the makeup choices were labeled as insufficient by MCNs not because of any clear platform rule, but because the MCN had already instilled the idea that visual performance is central to algorithmic success. As L22 put it, “They didn’t force me, but I felt like if I didn’t improve my makeup, I was being lazy or unprofessional.”
@@
MCNs’ interpretation is central to labor control because many streamers remain in the industry precisely through this form of hope-driven labor. Without this horizon of hope, many streamers would exit the industry much earlier. F1 worked as a streamer for one year, first collaborating with an MCN for six months and then continuing independently for another six months. During her time with the MCN, she strongly believed that success was attainable and that the algorithm rewarded effort in a fair manner. This belief persisted even after conflicts with the MCN emerged. As a result, she continued live-streaming independently for an additional six months, holding onto the expectation that consistent labor would eventually pay off. It was only after sustained poor performance and declining income that she decided to leave the industry. F1 explained: “Many of us stay in this industry because we believe there is hope to make money. MCNs always describe other industries as less fair than live-streaming.”
It is worth noting that most platforms, including Douyin, allow streamers to go live independently. Following the general principle, streamers do not need MCNs in order to participate in live-streaming. However, many streamers choose to join MCNs because they lack the necessary equipment, technical knowledge, and practical experience at the early stages of their careers. MCNs position themselves as providers of guidance, resources, and know-how, and their simplified narratives about algorithmic success make live-streaming appear worth sustained effort. These easily teachable interpretations are also central to MCNs’ ability to attract and retain a sufficiently large pool of streamers. Importantly, streamers are almost always free to leave. Especially after 2022, exit costs have become relatively low, with penalties described by MCNs as “basically one month’s salary” (M11). By framing success as something that can be achieved through effort, MCNs encourage streamers to join and to remain in the organization long enough to generate additional revenue streams, such as equipment rental fees discussed later in this paper.
Under these conditions, discursive construction becomes a key mechanism for labor control. Because streamers can exit at any time, control cannot rely on formal coercion. Instead, it depends on streamers’ willingness to stay, to keep investing labor, and to continue believing in the possibility of success.
@@ 5.2. Embedding Algorithmic Beliefs into Material and Contractual Structures
The interpretive work described in the previous section does not stop at motivation. MCNs extend their algorithmic explanations into the material, aesthetic, and contractual arrangements that define how streamers work. What begins as speculative beliefs, such as the idea that visual clarity or consistent appearance leads to higher algorithmic visibility, is gradually institutionalized into binding routines and financial commitments.
Some equipment upgrades can indeed improve technical conditions, such as network stability or image clarity. However, MCNs actively amplify the significance of these improvements by framing them as decisive factors for algorithmic recognition and success. Through this framing, incremental technical enhancements are transformed into ongoing requirements for investment.
@@
MCNs frequently tell streamers that clearer visuals and smoother audio will improve their ranking, reinforcing this belief through formal or informal expectations. L19 described his experience: “I spent 10,000 RMB on a Sony camera because the MCN said it would help me algorithmic visibility rank better. But nothing changed. They kept saying, maybe I need to improve my makeup or get a new background.”
@@
This expectation extends far beyond cameras. Ring lights, microphones, mixers, reflective backdrops, and LED panels are all framed as essential. In some cases, MCNs distribute internal “technical evaluation reports” that nudge streamers toward upgrades. M12 explained how this interpretation benefits the company: “Most streamers don’t buy equipment all at once. They rent from us first. If they start earning, they’ll invest more. If they don’t, at least we make money from the rentals.”
@@
Alongside hardware, MCNs standardize streamers’ aesthetic presentation. Guidelines often dictate not only tone and posture, but makeup, camera angles, and even clothing schedules. M9 described a typical protocol: “They’re not allowed to wear the same outfit in a two-week cycle. We even advise which colors look better on camera. Some of them rent clothes from us.”
@@
These requirements are enforced contractually through studio rules, performance evaluations, and surveillance systems. Even streamers who work from home are monitored. M4 explained: “We ask them to share their live link with a manager who checks if the lighting, camera, background are aligned with expectations. If not, they get a warning and we suggest them do a upgrade.”
@@
In some cases, MCNs also manage the physical layout of labor. Many operate shared studio spaces where streaming rooms are tiered by performance level. Higher-ranked streamers receive better infrastructure and services. L16 described how space and performance were linked: “At first, I shared a basement room with three other girls. The equipment frequently freezes or becomes unresponsive, it’s difficult to operate, and there’s poor air circulation. When I ranked higher one month, I moved to a private room upstairs. But I had to pay more for it, nearly 3,000 RMB a month.” These layered expectations create cycles of escalating cost and performance pressure. L20 recalled: “I spent most of my early earnings on room rental, clothes, gear. I didn’t even realize I was in debt until month three. But by then, I didn’t dare quit; I was too invested.” Streamers are often told that these are necessary “investments” in their future algorithmic visibility. But the platform algorithm offers no guarantees. L22 shared his experience: “I borrowed money from my cousin to redecorate my stream room. I told myself, if I looked more professional, I’d get better exposure. But nothing happened. Now I just feel stuck.” As shown in  Figure 2, the material differentiation of studio spaces reflects a managed hierarchy of visibility and professionalism, reinforcing MCNs’ externally imposed belief that algorithmic success is contingent on aesthetic and technical upgrades.
@@
As illustrated in subsubsection 4.2.1, MCNs build streamers’ belief that algorithmic success depends on visible performance cues and technical polish. M12 captured the disciplining effect of this algorithmic interpretation: “We don’t force anyone to stay. But when you’ve rented equipment, decorated your room, signed the platform contract; it’s hard to walk away.”
These behaviors and forms of self-understanding are shaped almost entirely by the algorithmic interpretations constructed and promoted by MCNs. In practice, live-streaming itself requires only minimal technical infrastructure, often no more than a smartphone with a built-in camera and microphone. Many platforms are basically functional under such basic conditions. The escalation of equipment purchases, studio upgrades, and aesthetic investments emerges only when streamers come to accept MCNs’ claims that algorithmic success depends on visible professionalism and constant upgrading. MCNs rarely compel streamers to make such purchases through direct coercion. Instead, investment is encouraged through interpretive frames that link spending to algorithmic recognition and future success.